Write an equation of the line in the form y =mx+b if the line has the same y-intercept as the line whose equation is 2y = 6x+8
and is parallel to the line whose equation is 4x + 4y = 20



Thanks Kat

Answer by ankor@dixie-net.com(22740)

Write an equation of the line in the form y = mx + b if the line has the same
y-intercept as the line whose equation is 2y = 6x+8
and is parallel to the line whose equation is 4x + 4y = 20
:
Find the y intercept from the equation 2y = 6x + 8
divide equation by 2
y = 3x + 4
y intercept = 4
:
Find the slope from the equation 4x + 4y = 20
divide equation by 4
x + y = 5
put in the slope/intercept form
y = -x + 5
m = -1 (parallel lines have the same slope)
:
The equation parallel to the above with a intercept of 4
y = -x + 4